OpenStudy (anonymous):

Evaluate x^2-4x+1 and x=2

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

substitute 2 for x

OpenStudy (anonymous):

plug 2 into the equation. so (2)^2 - 4(2) +1 = 4-8+1=-3

OpenStudy (anonymous):

-3